OpenStudy (anonymous):

find the largest no of 4 digits exactly divisible by 12,15,18,27.

OpenStudy (anonymous):

@apoorvk

OpenStudy (apoorvk):

Okay. so, find out the LCM of these nos. first of all. Now any no. that's divisible by all 4 of these nos., would be a multiple of that. So, find out the largest 4 digit multiple of the LCM!

OpenStudy (anonymous):

i got de lcm as54...den hw 2 do it??

OpenStudy (apoorvk):

Now which multiple of 54 is the largest 4 digit multiple possible? Find that by trial and error. 54 is a wee bit greater than 50, (and I know 50 times 200 is 10000, but I need a no. smaller than that). So, I check for 54 times 198, which is 9972, and is a 4 digit no. Now just confirming if this is the largest, I check for 19, but 54 times 19 = 10026, which is 5-digits. Hence, N/A. So, 54 times 18 it is, i.e. 9972!

OpenStudy (anonymous):

but how about 15. 9972/15=664.8!

OpenStudy (apoorvk):

That means your LCM was wrong. It is wrong in fact @prasannalakshmi :/ And i foolishly didn't even check *facepalm*

OpenStudy (anonymous):

oopz it was 540 infact

OpenStudy (anonymous):

the answer is 540*18 = 9720

OpenStudy (anonymous):

my english is not good, i wish my explain can help you. among 12, 15, 18, 27 the least number has the integer multiple of these is 540=3*3*3*4*5 and use the lesat 5 digits number 10000 devide it the answer would be 18.5...... so use the integer 18 times 540 which is 9720

OpenStudy (anonymous):

got it? :)

OpenStudy (anonymous):

ya:)

OpenStudy (anonymous):

thnq gyzz

OpenStudy (anonymous):

i dont understand your short write -.-

OpenStudy (apoorvk):

thnq = thank you gyzzz = guys don't go by the spelling, *feel* the pronunciation :P

OpenStudy (anonymous):

ya. it would be helpful to me to learn english!